# Stopping Rules for Stochastic Gradient Descent via Anytime-Valid Confidence Sequences

### v2

@@ Abstract
We study stopping rules for stochastic gradient descent (SGD) for convex optimization from the perspective of anytime-valid confidence sequences. Classical analyses of SGD provide convergence guarantees in expectation or at a fixed horizon, but offer no statistically valid way to assess, at an arbitrary time, how close the current iterate is to the optimum. We develop an anytime-valid, data-dependent upper confidence sequence for the weighted average suboptimality of projected SGD, constructed via nonnegative supermartingales and requiring no smoothness or strong convexity. This confidence sequence yields a simple stopping rule that is provably $\varepsilon$-optimal with probability at least $1-\alpha$ and is almost surely finite under standard stochastic approximation stepsizes. To the best of our knowledge, these are the first rigorous, time-uniform performance guarantees and finite-time $\varepsilon$-optimality certificates for projected SGD with general convex objectives, based solely on observable trajectory quantities.
@@ 1.2 Contributions
    Anytime-valid confidence sequence. Under Assumptions 1.1 and 1.2, Theorem 2.2 provides an upper confidence sequence $\{U_{t}(\alpha)\}_{t\geq 1}$ for the weighted average suboptimality $\bar{F}_{t}$ satisfying
@@
    The bound $U_{t}(\alpha)$ is fully data-dependent and remains valid uniformly over time, including under adaptive monitoring and data-dependent choices of stepsizes.
@@
    Anytime-valid $\varepsilon$-optimal stopping rule. Fix $\varepsilon>0$ and $\alpha\in(0,1)$. Theorem 3.1 proves that the stopping time $\tau_{\varepsilon}$ defined in (7) is almost surely finite under Assumption 1.2, and the returned iterate $\bar{x}_{\tau_{\varepsilon}}$ satisfies
@@
    Thus projected SGD, when equipped with the confidence sequence from (i), becomes an anytime-valid optimization procedure with a rigorous finite-time certificate of $\varepsilon$-optimality.
@@ 1.3 Related Work
Stochastic approximation and asymptotic normality. Stochastic approximation provides the theoretical foundation for stochastic gradient methods. The seminal paper [25] established almost sure convergence of iterative schemes under diminishing stepsizes, while [17] developed a comprehensive asymptotic theory based on ordinary differential equations, covering stability, asymptotic normality, and efficiency. A key advance showed that appropriately averaged iterates achieve asymptotically optimal covariance, providing a principled justification for studying weighted averages (which we adopt in this paper) of SGD trajectories [21]. This asymptotic perspective was sharpened in [9], which characterized minimax and asymptotically efficient rates for stochastic optimization, including constrained formulations, which was further extended to nonsmooth stochastic approximation in [6]. Together, these results motivate the use of averaged iterates and enable asymptotic confidence regions for optimal solutions based on limiting normality. However, such guarantees are inherently asymptotic: they do not yield finite-time confidence sets, nor do they provide time-uniform validity under continuous monitoring or adaptive stopping.
@@
Fixed-horizon high-probability guarantees for SGD. A vast literature develops nonasymptotic convergence guarantees for stochastic gradient (and mirror-descent-type) methods that hold with high probability at a pre-specified iteration horizon. Early robust frameworks for stochastic convex optimization established high-probability bounds under bounded or sub-Gaussian noise assumptions, providing optimal oracle complexity guarantees for stochastic approximation and stochastic programming [20, 18]. Parallel developments in machine learning emphasized refined nonasymptotic analyses of stochastic approximation algorithms, including explicit bias–variance tradeoffs and finite-time rates for averaged iterates [19, 2, 8]. Subsequent work sharpened fixed-horizon high-probability bounds across a range of regimes, including nonsmooth and strongly convex objectives, and clarified when optimal rates can be achieved by averaging versus last-iterate solutions [22, 27, 12, 13, 16]. Complementary information-theoretic lower bounds characterize the fundamental limits of stochastic convex optimization at fixed confidence levels [1]. More recent contributions extend fixed-horizon high-probability guarantees beyond light-tailed noise models, including heavy-tailed settings handled via gradient clipping and related robustification techniques [11, 3], as well as general boosting and robust-estimation wrappers that convert low-probability guarantees into high-confidence ones [5]. Together, these works provide sharp finite-time guarantees at prescribed horizons and confidence levels. However, these guarantees are intrinsically fixed-time: they do not yield confidence sets or optimality certificates that remain valid uniformly over time under continuous monitoring or adaptive stopping without additional union bounds or restarts, motivating the time-uniform approach pursued in this paper.
@@
Time-uniform inference and confidence sequences. Time-uniform (also called anytime-valid) inference provides uncertainty bounds that remain valid simultaneously over all times and are therefore robust to continuous monitoring and optional stopping. A foundational tool is Ville’s inequality for nonnegative supermartingales [28], which converts supermartingale constructions into time-uniform deviation guarantees. Early statistical instances of this idea include classical confidence sequences for basic parameters [4], as well as boundary-crossing results for martingales [26]. Modern developments systematize these principles through nonnegative supermartingale constructions and, in particular, mixture and stitching arguments, yielding sharp time-uniform concentration inequalities that adapt to variance-type processes [15]; see also related time-uniform Chernoff bounds based on exponential supermartingales [14] and the broader literature on self-normalized processes [7]. More recently, these ideas have been reframed through e-values and game-theoretic statistics, providing a unified perspective on optional-stopping-safe inference and a comprehensive entry point to the literature [24, 23]. While this framework has been developed primarily for sequential testing and estimation, it has begun to influence learning and decision-making problems, including anytime-valid inference for bandits and betting-based estimation [30, 29]. Our work leverages this time-uniform machinery in an optimization context, using supermartingale mixtures to construct suboptimality certificates that remain valid under arbitrary, data-dependent stopping.
@@
Closest related work. The works most closely related to ours study anytime or time-uniform guarantees for stochastic approximation, but differ substantially in assumptions and in the form of the resulting guarantees. Classical results such as [22] establish sharp high-probability convergence rates for SGD under strong convexity, bounded stochastic gradients, and fixed stepsizes, with guarantees stated for a pre-specified iteration horizon. As such, they do not support data-dependent stopping times. More recent work [10] derives anytime high-probability bounds for stochastic gradient methods with momentum under convexity and smoothness assumptions, using algorithm-specific dynamics and stepsize choices motivated by continuous-time ODE analysis. While these results control the optimization error uniformly over time, they do not provide explicit constants or trajectory-observable confidence sequences, and therefore cannot be used to derive certified stopping rules. In contrast to both works, our results apply to plain projected SGD under general convexity, yield fully explicit and trajectory-dependent confidence sequences for suboptimality, and lead to provably correct stopping rules that remain valid under arbitrary adaptive monitoring.
@@ 2 Anytime-Valid Confidence Sequences
and a trajectory-dependent quantity that governs the width of the confidence sequence: for each $t\geq 1$, define
$\sigma_{t}^{2}:=\sigma^{2}\eta_{t}^{2}Z_{t},\qquad\Sigma_{t}^{2}:=\sum_{s=1}^{t}\sigma_{s}^{2},$ (8)
where $\sigma^{2}>0$ is the sub-Gaussian variance proxy appearing in Assumption 1.1(iii). The cumulative term $\Sigma_{t}^{2}$ will emerge as a key scale parameter in our time-uniform concentration bound: larger values correspond to trajectories with larger stepsizes or larger distances to the optimum, and therefore lead to wider confidence sequences. Although $Z_{t}$ is not observable, it is always bounded by $R_{x}^{2}$, and keeping it explicit highlights which components of the trajectory influence the final bound. A fully observable, worst-case version can be obtained by replacing $Z_{t}$ with $R_{x}^{2}$.
@@
$U_{t}(\alpha):=\frac{1}{2S_{t}}\Bigg(C\,\sqrt{\Sigma_{t}^{2}\Big(\log\tfrac{1}{\alpha}+c+\log\log(\mathrm{e}+\Sigma_{t}^{2})\Big)}+Z_{1}+\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}\Bigg),$ (9)
where $C,c>0$ are constants, provided explicitly below. This boundary combines three sources of variability: the cumulative variance proxy $\Sigma_{t}^{2}$, the initial distance $Z_{1}$, and the accumulated squared gradient steps $\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}$. The leading term involving $\Sigma_{t}^{2}$ captures the stochastic fluctuations of the process and stems from a time-uniform concentration inequality applied to an adapted sequence satisfying conditional sub-Gaussian tail bounds (see the proof of Theorem 2.2). In contrast, the terms $Z_{1}$ and $\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}$ come directly from the deterministic structure of projected SGD, and appear when the distance recursion is telescoped.
Before proceeding, we introduce a mild technical condition on $\Sigma_{t}^{2}$ that simplifies our analysis.
Assumption 2.1 (Minimal cumulative variance proxy).
Fix $\alpha\in(0,0.6)$. We assume
$\Sigma_{t}^{2}\geq 2\Big(\log\frac{\pi^{2}}{6\alpha}+1\Big)\qquad\forall t\geq 1.$
This assumption is practically unrestrictive. Since $\Sigma_{t}^{2}=\sum_{s=1}^{t}\sigma^{2}\eta_{s}^{2}Z_{s}$, the quantity $\Sigma_{t}^{2}$ typically exceeds this threshold after only a few iterations; if not, one may simply slightly enlarge the first stepsize $\eta_{1}$. Moreover, we specify the constants appearing in $U_{t}(\alpha)$:
$C\;:=\;6\sqrt{\max\!\left\{1,\;\frac{\log\Big(\log\widetilde{\Sigma}_{\infty}^{2}+1\Big)}{\log\Big(\log(e+1)\Big)}\right\}},\qquad c\coloneqq\log\frac{\pi^{2}}{6}.$ (10)
Note that under Assumption 1.2, $\widetilde{\Sigma}_{\infty}^{2}:=\sigma^{2}R_{x}^{2}V_{\infty}<\infty$, with $V_{\infty}\coloneqq\sum_{s=1}^{\infty}\eta_{s}^{2}$, ensuring that $C$ is finite. With Assumption 2.1 and the definition of $C$, the confidence sequence $U_{t}(\alpha)$ in (9) is now fully specified. The next theorem establishes that $U_{t}(\alpha)$ forms an anytime-valid upper confidence sequence for the weighted suboptimality process $\{\bar{F}_{t}\}_{t\geq 1}$.
@@
Suppose Assumptions 1.1, 1.2, and 2.1 hold, and let $\sigma_{t}^{2}$ and $\Sigma_{t}^{2}$ be defined as in (8). Then, for every $\alpha\in(0,0.6)$, the process $\{U_{t}(\alpha)\}_{t\geq 1}$ defined in (9), with constant $C$ defined in (10), satisfies
@@
A noteworthy feature of the bound in Theorem 2.2 is its asymptotic decay. Under the stochastic approximation stepsize conditions $\sum_{t}\eta_{t}=\infty$ and $\sum_{t}\eta_{t}^{2}<\infty$, we have $S_{t}\to\infty$ and $\Sigma_{t}^{2}$ uniformly bounded, since the trajectory remains in a compact set and $\sigma_{t}^{2}=\sigma^{2}\eta_{t}^{2}Z_{t}$ is summable. Moreover, by Lemma A.1, the series $\sum_{s=1}^{\infty}\eta_{s}^{2}\|g_{s}\|^{2}$ converges almost surely to a finite random limit. Consequently, we have the following result.
Proposition 2.3 (Asymptotic decay of the confidence sequence).
Under the conditions of Theorem 2.2, for any fixed $\alpha\in(0,0.6)$,
$U_{t}(\alpha)\to 0\quad\text{almost surely as }t\to\infty.$
In particular, the confidence sequence is not only valid for all $t\geq 1$ but also shrinks to zero at the classical $1/S_{t}$ rate of stochastic approximation (up to a finite random constant).
@@
$\widetilde{\sigma}_{t}^{2}:=\sigma^{2}\eta_{t}^{2}R_{x}^{2},\qquad\widetilde{\Sigma}_{t}^{2}:=\sum_{s=1}^{t}\widetilde{\sigma}_{s}^{2}=\sigma^{2}R_{x}^{2}V_{t}.$
Then for every $\alpha\in(0,0.6)$ the process
$U_{t}^{\mathrm{obs}}(\alpha):=\frac{1}{2S_{t}}\Bigg(C\sqrt{\widetilde{\Sigma}_{t}^{2}\!\left(\log\tfrac{1}{\alpha}+c+\log\log(\mathrm{e}+\widetilde{\Sigma}_{t}^{2})\right)}+R_{x}^{2}+\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}\Bigg)$
@@
We highlight that for Corollary 2.5, the technical Assumption 2.1 need only be imposed on the observable variance proxy $\widetilde{\Sigma}_{t}^{2}$. Note that $U_{t}^{\mathrm{obs}}(\alpha)$ is $\mathcal{F}_{t}$-measurable and can be computed online at each iteration using only observable information.
@@
We now prove assertion (ii). From (16) and the convexity inequality just used, we also have the sharper pathwise bound
@@
Our goal is now to show that the partial sums $\bar{X}_{t}$ admit a time-uniform upper bound in terms of the cumulative variance proxy $\Sigma_{t}^{2}=\sum_{s=1}^{t}\sigma_{s}^{2}$. This is where the main probabilistic machinery enters: we construct an exponential supermartingale for each fixed $\lambda$, mix over a geometric grid of $\lambda$’s, and apply Ville’s inequality to obtain a high-probability bound that holds simultaneously for all $t$.
@@
$c_{1}\;:=\;\max\!\left\{1,\;\frac{\log\Big(\log\widetilde{\Sigma}_{\infty}^{2}+1\Big)}{\log\Big(\log(e+1)\Big)}\right\},$ (19)
with $\widetilde{\Sigma}_{\infty}^{2}:=\sigma^{2}R_{x}^{2}V_{\infty}<\infty$, and set $\lambda_{0}:=\sqrt{c_{1}/\sqrt{e}}$. Moreover, for $k=0,1,2,\ldots$, define
$\lambda_{k}:=\lambda_{0}e^{-k/2},\qquad w_{k}:=\frac{6}{\pi^{2}(k+1)^{2}}.$
Note that $\sum_{k\geq 0}w_{k}=1$. We are now ready to state the time-uniform concentration result for $\bar{X}_{t}$.
@@
Fix $\alpha\in(0,0.6)$, and let $\{X_{t}\}_{t\geq 0}$ be the process defined in (12). For $t\geq 0$, define the mixture process
$\bar{E}_{t}:=\sum_{k\geq 0}w_{k}\exp\!\left(\lambda_{k}\bar{X}_{t}-2\lambda_{k}^{2}\,\Sigma_{t}^{2}\right),$
@@
$\forall t\geq 1:\ \bar{X}_{t}\leq C\sqrt{\Sigma_{t}^{2}\Big(\log\frac{1}{\alpha}+c+\log\log(\mathrm{e}+\Sigma_{t}^{2})\Big)}.$
with constants $C,c$ defined in (10).
@@
Since this holds simultaneously for all $k\geq 0$, we get the deterministic bound
$\bar{X}_{t}\leq\inf_{k\geq 0}B_{k}(\Sigma_{t}^{2}).$ (20)
It remains to upper bound the infimum in terms of $\Sigma_{t}^{2}$ and $\alpha$.
Step (iii): Explicit form of $B_{k}(\Sigma_{t}^{2})$. For notational simplicity, set
$v:=\Sigma_{t}^{2}>0,\qquad L_{\alpha}:=\log\frac{\pi^{2}}{6\alpha}.$
Using $w_{k}={6}/(\pi^{2}(k+1)^{2})$ and $\lambda_{k}=\lambda_{0}e^{-k/2}$, we have
$\log\frac{1}{\alpha w_{k}}=\log\frac{\pi^{2}}{6\alpha}+2\log(k+1)=L_{\alpha}+2\log(k+1),$
@@
$\frac{1}{\lambda_{k}}=\frac{e^{k/2}}{\lambda_{0}}.$

### v3

@@ Abstract
We study stopping rules for stochastic gradient descent (SGD) for convex optimization from the perspective of anytime-valid confidence sequences. Classical analyses of SGD provide convergence guarantees in expectation or at a fixed horizon, but offer no statistically valid way to assess, at an arbitrary time, how close the current iterate is to the optimum. We develop an anytime-valid, data-dependent upper confidence sequence for the weighted average suboptimality of projected SGD, constructed via nonnegative supermartingales and requiring no smoothness or strong convexity. This confidence sequence yields a simple stopping rule that is provably $\varepsilon$-optimal with probability at least $1-\alpha$, with explicit bounds on the stopping time under standard stochastic approximation stepsizes. To the best of our knowledge, these are the first rigorous, time-uniform performance guarantees and finite-time $\varepsilon$-optimality certificates for projected SGD with general convex objectives, based solely on observable trajectory quantities.
@@ 1.2 Contributions
    Anytime-valid confidence sequence. Fix $\alpha\in(0,\tfrac{2}{\mathrm{e}})$. Theorem 2.1 provides an upper confidence sequence $\{U_{t}(\alpha)\}_{t\geq 1}$ for the weighted average suboptimality $\bar{F}_{t}$ satisfying
@@
    The bound $U_{t}(\alpha)$ is trajectory-adaptive and remains valid uniformly over time, including under adaptive monitoring and data-dependent choices of stepsizes. Moreover, a fully observable version of the confidence sequence, suitable for online implementation, is given in Corollary 2.4. Finally, under standard stochastic approximation stepsizes, $U_{t}(\alpha)$ decays at the rate $1/S_{t}$ almost surely; see Proposition 2.2 and Remark 2.3.
@@
    Anytime-valid $\varepsilon$-optimal stopping rule. Fix $\varepsilon>0$ and $\alpha\in(0,\tfrac{2}{\mathrm{e}})$. Theorem 3.1 proves that the stopping time $\tau_{\varepsilon}$ defined in (7) is almost surely finite, and the returned iterate $\bar{x}_{\tau_{\varepsilon}}$ satisfies
@@
    Thus projected SGD, when equipped with the confidence sequence from (i), becomes an anytime-valid optimization procedure with a rigorous finite-time certificate of $\varepsilon$-optimality. Explicit bounds on the stopping-time complexity for standard stepsize schedules are given in Propositions 3.4 and 3.5.
@@ 1.3 Related Work
Stochastic approximation and asymptotic normality. Stochastic approximation provides the theoretical foundation for stochastic gradient methods. The seminal paper [robbins1951stochastic] established almost sure convergence of iterative schemes under diminishing stepsizes, while [kushner2003stochastic] developed a comprehensive asymptotic theory based on ordinary differential equations, covering stability, asymptotic normality, and efficiency. A key advance showed that appropriately averaged iterates achieve asymptotically optimal covariance, providing a principled justification for studying weighted averages (which we adopt in this paper) of SGD trajectories [polyak1992acceleration]. This asymptotic perspective was sharpened in [duchi2021asymptotic], which characterized minimax and asymptotically efficient rates for stochastic optimization, including constrained formulations, which was further extended to nonsmooth stochastic approximation in [davis2024asymptotic]. Together, these results motivate the use of averaged iterates and enable asymptotic confidence regions for optimal solutions based on limiting normality. However, such guarantees are inherently asymptotic: they do not yield finite-time confidence sets, nor do they provide time-uniform validity under continuous monitoring or adaptive stopping.
@@
Fixed-horizon high-probability guarantees for SGD. A vast literature develops nonasymptotic convergence guarantees for stochastic gradient (and mirror-descent-type) methods that hold with high probability at a pre-specified iteration horizon. Early robust frameworks for stochastic convex optimization established high-probability bounds under bounded or sub-Gaussian noise assumptions, providing optimal oracle complexity guarantees for stochastic approximation and stochastic programming [nemirovski2009robust, lan2012optimal]. Related work also studied statistical validation of mirror-descent–type stochastic approximation methods, providing confidence bounds for the optimal objective value at a fixed iteration horizon rather than under adaptive or continuously monitored stopping rules [lan2012validation]. Parallel developments in machine learning emphasized refined nonasymptotic analyses of stochastic approximation algorithms, including explicit bias–variance tradeoffs and finite-time rates for averaged iterates [moulines2011non, bach2013non, dieuleveut2016nonparametric]. Subsequent work sharpened fixed-horizon high-probability bounds across a range of regimes, including nonsmooth and strongly convex objectives, and clarified when optimal rates can be achieved by averaging versus last-iterate solutions [rakhlin2012making, shamir2013stochastic, harvey2019tight, harvey2019simple, jain2021making]. Complementary information-theoretic lower bounds characterize the fundamental limits of stochastic convex optimization at fixed confidence levels [agarwal2012information]. More recent contributions extend fixed-horizon high-probability guarantees beyond light-tailed noise models, including heavy-tailed settings handled via gradient clipping and related robustification techniques [gorbunov2020stochastic, cutkosky2021high], as well as general boosting and robust-estimation wrappers that convert low-probability guarantees into high-confidence ones [davis2021low]. Together, these works provide sharp finite-time guarantees at prescribed horizons and confidence levels. However, these guarantees are intrinsically fixed-time: they do not yield confidence sets or optimality certificates that remain valid uniformly over time under continuous monitoring or adaptive stopping without additional union bounds or restarts, motivating the time-uniform approach pursued in this paper. Viewed differently, the present work does not seek to refine convergence rates for SGD, but rather to introduce a complementary, time-uniform notion of statistical validity for stochastic optimization algorithms, in which stopping times are treated as random, data-dependent objects subject to formal inference.
@@
Time-uniform inference and confidence sequences. Time-uniform (also called anytime-valid) inference provides uncertainty bounds that remain valid simultaneously over all times and are therefore robust to continuous monitoring and optional stopping. A foundational tool is Ville’s inequality for nonnegative supermartingales [ville1939], which converts supermartingale constructions into time-uniform deviation guarantees. Early statistical instances of this idea include classical confidence sequences for basic parameters [darling1967confidence], as well as boundary-crossing results for martingales [robbins1970boundary]. Modern developments systematize these principles through nonnegative supermartingale constructions and, in particular, mixture and stitching arguments, yielding sharp time-uniform concentration inequalities that adapt to variance-type processes [howard2021time]; see also related time-uniform Chernoff bounds based on exponential supermartingales [howard2020time] and the broader literature on self-normalized processes [de2009self]. More recently, these ideas have been reframed through e-values and game-theoretic statistics, providing a unified perspective on optional-stopping-safe inference and a comprehensive entry point to the literature [ramdas2023game, ramdas2025hypothesis]. While this framework has been developed primarily for sequential testing and estimation, it has begun to influence learning and decision-making problems, including anytime-valid inference for bandits and betting-based estimation [waudby2024anytime, waudby2024estimating]. Our work leverages this time-uniform machinery in an optimization context, using supermartingale mixtures to construct suboptimality certificates that remain valid under arbitrary, data-dependent stopping.
@@
Closest related work. The works most closely related to ours study anytime or time-uniform guarantees for stochastic approximation, but differ substantially in assumptions and in the form of the resulting guarantees. Classical results such as [rakhlin2012making] establish high-probability convergence bounds for SGD under strong convexity and bounded stochastic gradients, with guarantees stated uniformly over iterations up to a pre-specified horizon. As such, they do not support data-dependent stopping times. More recent work [feng2023anytime] derives anytime high-probability bounds for stochastic gradient methods with momentum under convexity and smoothness assumptions, using algorithm-specific dynamics and stepsize choices motivated by continuous-time ODE analysis. While these results control the optimization error uniformly over time, they do not provide explicit constants or trajectory-observable confidence sequences, and therefore cannot be used to derive certified stopping rules. Finally, [yu2025root] studies one-dimensional stochastic root finding and proposes a bisection-based procedure that yields time-uniform confidence intervals for the unknown root, but their guarantees concern inference on the unknown root rather than certificates of optimization suboptimality for projected SGD. In contrast to these works, our results apply to plain projected SGD under general convexity, yield fully explicit and trajectory-dependent confidence sequences for suboptimality, and lead to provably correct stopping rules that remain valid under arbitrary adaptive monitoring.
@@ 2 Anytime-Valid Confidence Sequences
and a trajectory-dependent quantity that enters the width of the confidence sequence: for each $t\geq 1$, define
$\sigma_{t}^{2}:=\sigma^{2}\eta_{t}^{2}Z_{t},\qquad\Sigma_{t}^{2}:=\sum_{s=1}^{t}\sigma_{s}^{2},\qquad\Sigma_{t,\mathrm{eff}}^{2}:=\max\!\left\{\Sigma_{t}^{2},\;2\Big(\log\tfrac{2}{\alpha}+1\Big)\right\}.$ (8)
where $\sigma^{2}>0$ is the sub-Gaussian variance proxy appearing in Assumption 1.1(iii). The quantity $\Sigma_{t,\mathrm{eff}}^{2}$ will be referred to as the effective cumulative variance proxy. It coincides with the true cumulative variance proxy $\Sigma_{t}^{2}$ once the latter exceeds a fixed threshold, and serves to prevent small-variance degeneracies in the time-uniform concentration argument. Larger values of $\Sigma_{t,\mathrm{eff}}^{2}$ correspond to trajectories with larger stepsizes or larger distances to the optimum, and therefore lead to wider confidence sequences. Although $Z_{t}$ is not directly observable, it is always bounded by $R_{x}^{2}$, and keeping it explicit highlights which components of the trajectory influence the final bound. A fully observable version of the confidence sequence is provided in Corollary 2.4 by replacing $Z_{t}$ with $R_{x}^{2}$.
@@
$U_{t}(\alpha):=\frac{1}{2S_{t}}\Bigg(C\,\sqrt{\Sigma_{t,\mathrm{eff}}^{2}\Big(\log\tfrac{2}{\alpha}+\log\log(\mathrm{e}+\Sigma_{t,\mathrm{eff}}^{2})\Big)}+Z_{1}+\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}\Bigg),$ (9)
where $C=C(\alpha)>0$ is an explicit constant, provided below. This boundary combines three sources of variability: the effective cumulative variance proxy $\Sigma_{t,\mathrm{eff}}^{2}$, the initial distance $Z_{1}$, and the accumulated squared gradient steps $\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}$. The leading term involving $\Sigma_{t,\mathrm{eff}}^{2}$ captures the stochastic fluctuations of the process and stems from a time-uniform concentration inequality applied to an adapted sequence satisfying conditional sub-Gaussian tail bounds (see the proof of Theorem 2.1). In contrast, the terms $Z_{1}$ and $\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}$ come directly from the deterministic structure of projected SGD, and appear when the distance recursion is telescoped. Before proceeding, we specify the constant $C$ appearing in $U_{t}(\alpha)$:
$C\;:=\;6\sqrt{\max\!\left\{1,\;\frac{\log\Big(\log\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}+1\Big)}{\log\Big(\log(\mathrm{e}+1)\Big)}\right\}},$ (10)
with $\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}$ defined as follows
$V_{\infty}\coloneqq\sum_{s=1}^{\infty}\eta_{s}^{2},\qquad\widetilde{\Sigma}_{\infty}^{2}:=\sigma^{2}R_{x}^{2}V_{\infty},\qquad\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}:=\max\!\left\{\widetilde{\Sigma}_{\infty}^{2},\;2\Big(\log\tfrac{2}{\alpha}+1\Big)\right\}.$ (11)
Since $V_{\infty}<\infty$, it follows that $\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}<\infty$, ensuring that the constant $C$ is finite. With the definition of $C$, the confidence sequence $U_{t}(\alpha)$ in (9) is now fully specified. The next theorem establishes that $U_{t}(\alpha)$ forms an anytime-valid upper confidence sequence for the weighted suboptimality process $\{\bar{F}_{t}\}_{t\geq 1}$.
@@
Suppose Assumptions 1.1 and 1.2 hold. Then, for every $\alpha\in(0,\tfrac{2}{\mathrm{e}})$, the process $\{U_{t}(\alpha)\}_{t\geq 1}$ defined in (9), with constant $C$ defined in (10), satisfies
@@
A noteworthy feature of the bound in Theorem 2.1 is its rate-controlled decay. Under the stochastic approximation stepsize conditions in Assumption 1.2, we have $S_{t}\to\infty$ and $\Sigma_{t,\mathrm{eff}}^{2}$ uniformly bounded, since the trajectory remains in a compact set and $\sigma_{t}^{2}=\sigma^{2}\eta_{t}^{2}Z_{t}$ is summable. Moreover, by Lemma A.1, the partial sums $\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}$ converge almost surely to a finite random limit. Consequently, we have the following result.
Proposition 2.2 (Rate-controlled almost-sure decay of the confidence sequence).
Under the conditions of Theorem 2.1, for any fixed $\alpha\in(0,\tfrac{2}{\mathrm{e}})$,
$U_{t}(\alpha)=O\!\left(\frac{1}{S_{t}}\right)\quad\text{almost surely},$
and in particular $U_{t}(\alpha)\to 0$ almost surely as $t\to\infty$.
@@
$\widetilde{\sigma}_{t}^{2}:=\sigma^{2}\eta_{t}^{2}R_{x}^{2},\qquad\widetilde{\Sigma}_{t}^{2}:=\sum_{s=1}^{t}\widetilde{\sigma}_{s}^{2}=\sigma^{2}R_{x}^{2}V_{t},\qquad\widetilde{\Sigma}_{t,\mathrm{eff}}^{2}:=\max\!\left\{\widetilde{\Sigma}_{t}^{2},\;2\Big(\log\tfrac{2}{\alpha}+1\Big)\right\}.$
Then for every $\alpha\in(0,\tfrac{2}{\mathrm{e}})$ the process
$U_{t}^{\mathrm{obs}}(\alpha):=\frac{1}{2S_{t}}\Bigg(C\sqrt{\widetilde{\Sigma}_{t,\mathrm{eff}}^{2}\!\left(\log\tfrac{2}{\alpha}+\log\log(\mathrm{e}+\widetilde{\Sigma}_{t,\mathrm{eff}}^{2})\right)}+R_{x}^{2}+\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}\Bigg)$
@@
Note that $U_{t}^{\mathrm{obs}}(\alpha)$ is $\mathcal{F}_{t}$-measurable and can be computed online at each iteration using only observable information.
@@
We now prove assertion (ii). From (17) and the convexity inequality just used, we also have the pathwise bound
@@
Our goal is now to show that the partial sums $\bar{X}_{t}$ admit a time-uniform upper bound in terms of the effective cumulative variance proxy $\Sigma_{t,\mathrm{eff}}^{2}$ defined in (8). This is where the main probabilistic machinery enters: we construct an exponential supermartingale for each fixed $\lambda$, mix over a geometric grid of $\lambda$’s, and apply Ville’s inequality to obtain a high-probability bound that holds simultaneously for all $t$.
@@
$c_{1}\;:=\;\max\!\left\{1,\;\frac{\log\Big(\log\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}+1\Big)}{\log\Big(\log(\mathrm{e}+1)\Big)}\right\},$ (20)
with $\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}$ defined in (11), and set $\lambda_{0}:=\sqrt{c_{1}/\sqrt{\mathrm{e}}}$. Moreover, for $k=0,1,2,\ldots$, define
$\lambda_{k}:=\lambda_{0}\mathrm{e}^{-k/2},\qquad w_{k}:=\frac{6}{\pi^{2}(k+1)^{2}}.$
Note that $\sum_{k=0}^{\infty}w_{k}=1$. We are now ready to state the time-uniform concentration result for $\bar{X}_{t}$.
@@
Fix $\alpha\in(0,\tfrac{2}{\mathrm{e}})$, and let $\{X_{t}\}_{t\geq 0}$ be the process defined in (13). For $t\geq 0$, define the mixture process
$\bar{E}_{t}:=\sum_{k=0}^{\infty}w_{k}\exp\!\left(\lambda_{k}\bar{X}_{t}-2\lambda_{k}^{2}\,\Sigma_{t}^{2}\right),$
@@
$\forall t\geq 1:\ \bar{X}_{t}\leq C\sqrt{\Sigma_{t,\mathrm{eff}}^{2}\Big(\log\frac{2}{\alpha}+\log\log(\mathrm{e}+\Sigma_{t,\mathrm{eff}}^{2})\Big)}.$
with constant $C=6\sqrt{c_{1}}$ and $c_{1}$ defined in (20).
@@
Noticing that $B_{k}(\cdot)$ is nondecreasing and $\Sigma_{t}^{2}\leq\Sigma_{t,\mathrm{eff}}^{2}$, we further have
$\bar{X}_{t}<B_{k}(\Sigma_{t,\mathrm{eff}}^{2})\qquad\forall k\geq 0.$
and since this holds simultaneously for all $k\geq 0$, we obtain the bound
$\bar{X}_{t}\leq\inf_{k\geq 0}B_{k}(\Sigma_{t,\mathrm{eff}}^{2}).$ (21)
It remains to upper bound the infimum in terms of $\Sigma_{t,\mathrm{eff}}^{2}$ and $\alpha$.
Step (iii): Explicit form of $B_{k}(\Sigma_{t,\mathrm{eff}}^{2})$. For notational simplicity, set
$v:=\Sigma_{t,\mathrm{eff}}^{2},\qquad L_{\alpha}:=\log\frac{2}{\alpha}\geq\log\frac{\pi^{2}}{6\alpha}.$
Using $w_{k}={6}/(\pi^{2}(k+1)^{2})$ and $\lambda_{k}=\lambda_{0}\mathrm{e}^{-k/2}$, we have
$\log\frac{1}{\alpha w_{k}}=\log\frac{\pi^{2}}{6\alpha}+2\log(k+1)\leq L_{\alpha}+2\log(k+1),$
@@
$\frac{1}{\lambda_{k}}=\frac{\mathrm{e}^{k/2}}{\lambda_{0}}.$